OpenStudy (anonymous):

Teach me how to solve this 3x^2-2x=8

OpenStudy (anonymous):

how to put fraction inside of square root?

OpenStudy (anonymous):

3x^2-6x+4x-8=0 3x(x-2)+4(x-2)=0 (3x+4)(x-2)=0 x=-4/3,2

OpenStudy (anonymous):

3x^2-2x=8 9xsquared-2x=8 -8 -8 (3x-4)(3x+2)=0 x=-4/3,2

OpenStudy (anonymous):

lol

terenzreignz (terenzreignz):

@CUTE_AKKI Quadratic formula, if you will :) @BoomShakalaka163 messaged me to use that method specifically

terenzreignz (terenzreignz):

Very well then, @BoomShakalaka163 Show your progress.

OpenStudy (anonymous):

how can i put fraction inside of square root?

terenzreignz (terenzreignz):

using LaTeX?

terenzreignz (terenzreignz):

just draw it if it's easier for you :)

OpenStudy (anonymous):

3x^-2x=8 ~> 3x^2-2x-8=0 a=3 b=2=c8 x=-(2)\[\frac{ -2 \pm \sqrt{(-2^{2})}-4(3)(8) }{ 2(3) }\] \[\frac{ 2 \pm \sqrt{4-96} }{6 }\] \[\frac{ 2+92 }{6 }\] but i dont think 94 is divisible by 6

terenzreignz (terenzreignz):

c = -8 just sayin'

OpenStudy (anonymous):

yes i transpose it 3x^-2x=8 ~> 3x^2-2x-8=0

terenzreignz (terenzreignz):

Yes... that's why c = -8 but you used c = 8

OpenStudy (anonymous):

oh yes, typo sorry. can you help me finish this.

terenzreignz (terenzreignz):

Yeah... redo your solution, but use c = -8, and show me what happens.

OpenStudy (anonymous):

it will be just -96

terenzreignz (terenzreignz):

and since there was a minus sign, it will become + 96, correct?

OpenStudy (anonymous):

so it will be 4+96?

terenzreignz (terenzreignz):

yup :)

terenzreignz (terenzreignz):

So it becomes... \[\Large \frac{2+\pm \sqrt{4+96}}{4}\]

OpenStudy (anonymous):

why 4? 2(3) is 6 am i right

terenzreignz (terenzreignz):

typo :) Sorry... replace with 6 :)

OpenStudy (anonymous):

so x1 2+100/6=102/6 =17 and x2 2-100/6 = 98/6 = ? its invalid

terenzreignz (terenzreignz):

don't forget that 4+96 was under a square-root sign...

OpenStudy (anonymous):

it will become 25. 2+100/25 = 5 2-100/25= ..

terenzreignz (terenzreignz):

\[\Large \frac{2+\pm \color{red}{\sqrt{4+96}}}{6}=\frac{2\pm \sqrt{100}}{6}\] from here...?

OpenStudy (anonymous):

doesnt it get removed after it was added? the sqt.

terenzreignz (terenzreignz):

Square root of 100, lol, what is it?

OpenStudy (anonymous):

10

terenzreignz (terenzreignz):

Okay, so it becomes... \[\Large \frac{2\pm10}{6}\]

OpenStudy (anonymous):

but still if 2+10/6 =12/6=2 2-10/6 =8/6 =??

terenzreignz (terenzreignz):

it's not 8/6 but -8/6 And it's equal to -4/3, simplified, which, if you notice, is one of the (correct) answers of @CUTE_AKKI

OpenStudy (anonymous):

ok so x1 = 2 x2 = -4/3?

OpenStudy (anonymous):

am i correct with this guys? x^2+8x-9 = 0 (x+9)(x-1)=0 x+9=0 x-1=0 x=9 x=-1

terenzreignz (terenzreignz):

Nope... It's almost correct... but if (x+9) = 0 doesn't that mean x = -9 and not x = 9? think about it :)

OpenStudy (anonymous):

i used quadratic equation by factoring

terenzreignz (terenzreignz):

Yeah... and you factored well, it's the step that followed that was faulty.

OpenStudy (anonymous):

btw im stuck at something let me show you. 8x^2+14x=-5 8x+14x+5=0 (2x+1)(4x+5) 2x+1=0 4x+5=0 x=2+1? x=4+5?

OpenStudy (anonymous):

ok then x =-9 x=1 too.

terenzreignz (terenzreignz):

\[2x+1 = 0\]\[2x = -1\]\[x = -\frac12\]nice and easy :)

OpenStudy (anonymous):

4x+5=0 4x=-5 x=5/4

terenzreignz (terenzreignz):

Don't be careless with signs. It's -5/4

OpenStudy (anonymous):

are you a student also?

terenzreignz (terenzreignz):

Yup :)

OpenStudy (anonymous):

you helped me almost everyday, but i always change in different usernames. lol

terenzreignz (terenzreignz):

Who goes there?

OpenStudy (anonymous):

yesterday i was little help same as the other day. i cant verify my email thats why

terenzreignz (terenzreignz):

Well... just here playing around.. :P

OpenStudy (anonymous):

so can you help me with my last question its about completing the square x^2-3x-4=0?

terenzreignz (terenzreignz):

Factor.

terenzreignz (terenzreignz):

oh... I really should learn to read... sorry... hang on.

OpenStudy (anonymous):

(x-4)(x+1)

OpenStudy (anonymous):

i think its not the correct step

terenzreignz (terenzreignz):

Yeah, that's not necessary,but since you already did it, I guess you already know the answers?

OpenStudy (anonymous):

this was the given example \[x-12=4x \] \[x^2-4x=12\] \[x^2-4x+4=12+4\]\[\sqrt{(x-2)^2=16}\]\[x-2=\pm4\] x1=4+2=6 x2=-4+x=-2

terenzreignz (terenzreignz):

Yes, I know how to complete the square (thank goodness) but since you already factored it into (x-4)(x+1) = 0 you already know the answers? I'm just doing this so that you know that we can also get to the correct answer using completing the square.

OpenStudy (anonymous):

nope not yet.

terenzreignz (terenzreignz):

Well, much like earlier when you solved by factoring... remember when you ahd (x+9)(x-1)=0 the solutions were x = -9 and x = 1 ? Same thing basically.

OpenStudy (anonymous):

btw am i right on this 7x^2+19x=6 7x^2+19x-6=0 (7x-3)(x+2) 7x-3=0 x+2=0 x=3/7 x=-2

terenzreignz (terenzreignz):

Yes.

OpenStudy (anonymous):

i dont get the x2−4x+4=12+4

terenzreignz (terenzreignz):

Okay... standard form for "completing the square" is this... \[\LARGE x^2 +\color{red}bx = \color{green}c\]

terenzreignz (terenzreignz):

Catch me so far?

OpenStudy (anonymous):

so x^2-3x-4=0 x^2-3x=4 x^2 and then im lost

terenzreignz (terenzreignz):

Yeah... relax...deep breaths... just look at that general form I gave you...

terenzreignz (terenzreignz):

Now, the first step is to consider b... the number next to the x (with no exponent...) In this case x^2 - 3x - 4 What is b?

OpenStudy (anonymous):

b is 3

terenzreignz (terenzreignz):

Okay, so what I want you to do with it... is to take half of it... what's half of 3?

OpenStudy (anonymous):

1.5

terenzreignz (terenzreignz):

yes. or 3/2 I think you should use 3/2... Now, square that number, you get?

OpenStudy (anonymous):

x^2-3x+3=4+3?

terenzreignz (terenzreignz):

Not 3... I said square the number, not double it... there's a difference :P

OpenStudy (anonymous):

\[\sqrt{\left( x-3^2\right) =7}\]

terenzreignz (terenzreignz):

You're getting WAY ahead of yourself, and you're doing it wrong...

OpenStudy (anonymous):

x^2 - 3x +3^2- 4

terenzreignz (terenzreignz):

No... it's not 3 you square, but 3/2.

OpenStudy (anonymous):

i cant seem to understand

terenzreignz (terenzreignz):

Okay... example... x² - 6x - 16 = 0 x² - 6x = 16 Get half of 6, which is 3, and then square it... getting 9. Add that to both sides... x² - 6x + 9 = 16 + 9 (x - 3)² = 25 Get the square root of both sides x - 3 = +- 5 x = 3 +- 5 x = 8 or -2

OpenStudy (anonymous):

i now know the steps but i dont know what to do in 3/2

terenzreignz (terenzreignz):

square it... you get 9/4 :)

terenzreignz (terenzreignz):

and add that to both sides, giving you... \[\Large x^2 -3x + \frac94 = 4 + \frac94\] \[\Large\left( x- \frac32\right)^2 = \frac{25}4\] solve from here.

OpenStudy (anonymous):

x-3/2^2=25/4 x-3/2=+-5/2 x-3/2=5/2 x1=5/2+3/2=8/4 =4/2 = 2 x2=-5/2+3/2 = 2/2 = 1 am i correct?

terenzreignz (terenzreignz):

How come 5/2 + 3/2 becomes 8/4 Did you add denominators? As for -5/2 + 3/2 = 2/2 I told you once, I told you a thousand times... PAY ATTENTION TO SIGNS.

OpenStudy (anonymous):

yes i did add.

OpenStudy (anonymous):

sorry, i get very forgetful.

terenzreignz (terenzreignz):

You don't add denominators. Big No-no

OpenStudy (anonymous):

oh, sorry so its 8/2 = 4 -2/2 = 1

terenzreignz (terenzreignz):

Seriously o.O Watch the signs...

terenzreignz (terenzreignz):

Or else review multiplication and division of integers.

OpenStudy (anonymous):

i will do that so am i right? 8/2 = 4 -2/2 = -1

terenzreignz (terenzreignz):

Okay, better.

OpenStudy (anonymous):

thanks again for clearing me that terrenz. but im still lost at the quadratic formula. x1=2 x2=-4/3?

terenzreignz (terenzreignz):

Yes, those are the answers.

OpenStudy (anonymous):

ok thanks, ill try to review it now and clear it more enough lol, btw is the factor of 4x^2-36= (4x-4)(x-9)?

terenzreignz (terenzreignz):

No.

OpenStudy (anonymous):

4(x+3)(x-3)?

terenzreignz (terenzreignz):

That's better.